A gas has a pressure of 5.25 atm and a volume of 7095 mL at 57°C. How many moles are in the sample? Use R=0.0821 atm L/mol K. 
Answer #1

Given:

P = 5.25 atm

V = 7095.0 mL

= (7095.0/1000) L

= 7.095 L

T = 57.0 oC

= (57.0+273) K

= 330 K

find number of moles using:

P * V = n*R*T

5.25 atm * 7.095 L = n * 0.0821 atm.L/mol.K * 330 K

n = 1.37 mol

Answer: 1.37 mol